Problem 3. Choose one answer: Due to plae buildup, a small part of a patients aorta has a smaller radius than a regular, healthy aorta. Which of heollowing is true regarding the unhealthy and healthy aorta? (a) Flow in the clogged part of the unhealthy aorta will have a lower velocity and higher pressure than in a healthy aorta. (b) Flow in the clogged part of the unhealthy aorta will have a higher velocity and higher pressure than in a healthy aorta (c) Flow in the clogged part of the unhealthy rtawl have a lower velocity and lower pressure than in a healthy aorta d) Flow in the clogged part of the unhealthy aorta will have a higher velocity and lowe r pressure than in a healthy aorta. (e) There w be no difference in the blood flow in the two cases.
Answer #1

The flow has to satisfy the continuity equation as mass remains conserved.

From continuity, we have A1*V1 = A2*V2

Hence velocity is inversely proportional to the area of cross section

In the clogged area velocity has to increase as area reduces.

For pressure, we can use the Bernoulli principle (assuming flow is incompressible and inviscid),

i,e. Energy remains conserved as P +0.5*RHO*V^2 = E RHO is the density of blood,

Hence pressure has to decrease as velocity increases.

Hence the answer is (d)